# Phonon hydrodynamics in bulk insulators and semi-metals

### 21 Feb 2024 v1

@@ Abstract
Decades ago, Gurzhi proposed that if momentum-conserving collisions prevail among heat-carrying phonons in insulators and charge-carrying electrons in metals, hydrodynamic features will become detectable. In this paper, we will review the experimental evidence emerging in the last few years supporting this viewpoint and raising new questions. The focus of the paper will be bulk crystals without (or with a very dilute concentration) of mobile electrons and steady-state thermal transport. We will also discuss the possible link between this field of investigation and other phenomena, such as the hybridization of phonon modes and the phonon thermal Hall effect.
@@ I Introduction
In his seminal 1968 paper, entitled “Hydrodynamic effects in solids at low temperatures”[1], Gurzhi wrote: “The phenomena of thermal conductivity of dielectric [i.e. insulators] and the electrical conductivity of metals have specific properties. In both cases, the total quasi-particle current turns out to be non-vanishing. It follows that when only normal collisions occur in the system, there could exist an undamped current in the absence of an external field which could sustain it.” The adjective ‘normal’ is employed here to specify those collisions (we will call them ‘Normal’), which are not of the Umklapp type, i.e. between two quasi-particles with wave-vectors whose sum exceeds the width of the Brillouin zone.
@@
Now, there is a fundamental difference between electrons and phonons regarding the Umklapp collisions. The wave-vector of a thermally excited phonon shrinks with temperature. At a given temperature, it is , where is the sound velocity, is the Boltzmann constant, and is the reduced Planck constant. Therefore, in order to perform Gurzhi’s thought experiment for phonons, all you need [to get rid of Umklapp events] is to cool down your sample well below a temperature, where , which corresponds to a fraction of the Debye temperature. The situation is quite different for electrons, whose wavevector is set by the Fermi radius and does not change with temperature. A very small Fermi surface, located at the center of the Brillouin zone [3] is needed to avoid Umklapp collisions.
@@
These fundamental differences may explain, at least partially, the historical fact that phonon hydrodynamics was explored before electron hydrodynamics. As early as 1974, two decades before the first recorded experimental investigation of electron hydrodynamics [4], Beck and co-workers [5] wrote a review paper on phonon hydrodynamics. However, after a promising start, research on phonon hydrodynamics went through hibernation in the last decades of the previous century.
The renaissance of this field of research began in 2015 with theoretical studies of heat transport in graphene [6, 7]. The present paper summarizes the evidence gathered in the last few years by steady-sate transport measurements, confirming the existence of a hydrodynamic window for thermal transport by phonons. Our main focus will be a subset of such hydrodynamic phenomena associated with the Poiseuille regime observed in bulk crystals by steady-state heat transport studies. For a recent review paper on phonon hydrodynamics, written from a theoretical perspective, and covering a broader range of topics, see [8].

### 08 Apr 2024 v2

@@ Abstract
Decades ago, Gurzhi proposed that if momentum-conserving collisions prevail among heat-carrying phonons in insulators and charge-carrying electrons in metals, hydrodynamic features will become detectable. In this paper, we will review the experimental evidence emerging in the last few years supporting this viewpoint and raising new questions. The focus of the paper will be bulk crystals without (or with a very dilute concentration of) mobile electrons and steady-state thermal transport. We also discuss the possible link between this field of investigation and other phenomena, such as the hybridization of phonon modes and the phonon thermal Hall effect.
@@ I Introduction
In his seminal 1968 paper, entitled “Hydrodynamic effects in solids at low temperatures”[1], Gurzhi wrote: “The phenomena of thermal conductivity of dielectric [i.e. insulators] and the electrical conductivity of metals have specific properties. In both cases, the total quasi-particle current turns out to be non-vanishing. It follows that when only normal collisions occur in the system, there could exist an undamped current in the absence of an external field which could sustain it.” The adjective ‘normal’ is employed here to specify those collisions (we will call them ‘Normal’), which are not of the Umklapp type, i.e. do not produce a change in quasimomentum larger than the width of the Brillouin zone.
@@
Now, there is a fundamental difference between electrons and phonons regarding the Umklapp collisions. The wave-vector of a thermally excited phonon decreases with temperature. At a given temperature, it is , where is the sound velocity, is the Boltzmann constant, and is the reduced Planck constant. Therefore, in order to perform Gurzhi’s thought experiment for phonons, all you need [to get rid of Umklapp events] is to cool down your sample well below a temperature, where , which corresponds to a fraction of the Debye temperature. The situation is quite different for electrons, whose wavevector is set by the Fermi radius and does not change with temperature. A very small Fermi surface, located at the center of the Brillouin zone [3] is needed to avoid Umklapp collisions.
@@
These fundamental differences may explain, at least partially, the historical fact that phonon hydrodynamics was explored before electron hydrodynamics. As early as 1974, two decades before the first recorded experimental investigation of electron hydrodynamics [4], Beck and co-workers [5] wrote a review paper on phonon hydrodynamics.
Second sound refers to the wave-like propagation of temperature and entropy [6]. It was first observed in superfluid helium by Peshkov in 1946 [7, 8]. In 1952, Dingle suggested that a density fluctuation can propagate as a second sound wave in a phonon gas [9]. Guyer and Krumhansl [10] argued that the Poiseuille flow of phonons in a steady-state experiment is concomitant with the occurrence of phononic second sound, both requiring the same hierarchy of scattering times. Thanks to several works [11, 12, 13, 14], it was established that a hydrodynamic window emerges between the ballistic and diffusive transport regimes when momentum-conserving collisions between phonons become sufficiently abundant.
After a promising start, this field of research went through hibernation in the last decades of the previous century. The renaissance of this field of research began in 2015 with theoretical studies of heat transport in graphene [15, 16]. The present paper summarizes the evidence gathered in the last few years by steady-sate transport measurements, confirming the existence of a hydrodynamic window for thermal transport by phonons. Our main focus will be a subset of such hydrodynamic phenomena associated with the Poiseuille regime observed in bulk crystals by steady-state heat transport studies. For a recent review paper on phonon hydrodynamics, written from a theoretical perspective, and covering a broader range of topics, see [17].